# Conditional Wasserstein Distances with Applications in Bayesian OT Flow Matching

### v2

@@ 5 Geodesics and velocity fields
In this section, we deal with geodesics and velocity fields in $\left(\mathcal{P}_{Y,2}(\mathbb{R}^{d}\times\mathbb{R}^{m}),W_{2,Y}\right)$. We restrict our attention to $p=2$ and $A=\mathbb{R}^{d}$, $B=\mathbb{R}^{m}$. Coming from inverse problems, we have considered probability measures $P_{X,Y}$ related to random variables $(Y,X)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$. When switching to flows, it is more convenient to address equivalently just probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{m}$.
Let us recall some results which can be found, e.g. in [6] for our setting. A curve $\mu\colon[0,1]\to\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ is called a geodesic if
@@
The Wasserstein space is geodesic, i.e. any two measures $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ can be connected by a geodesic. Let $e_{t}:(\mathbb{R}^{d}\times\mathbb{R}^{m})^{2}\to\mathbb{R}^{n}\times\mathbb{R}^{m}$, $t\in[0,1]$ by defined by
@@
Conversely, any optimal plan $\alpha\in\Gamma(\mu_{0},\mu_{1})$ gives by (16) rise to a geodesic connecting $\mu_{0}$ and $\mu_{1}$. The following lemma considers curves defined by (16) which connect measures $\mu_{0},\mu_{1}\in\mathcal{P}_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$. Its proof is given in the appendix and is similar to [6, Theorem 7.2.2].
Lemma 5.
@@
By the following proposition, the above geodesic $\mu_{t}$ has an associated vector field $v_{t}$ such that $(\mu_{t},v_{t})$ satisfy a continuity equation. Moreover, informally speaking, the associated vector field $v_{t}$ does not transport any mass in the $y$-component. This is related to the observation in [20, Section 4.2].
Proposition 6.
@@
The proof is given in Appendix D and parts $i),ii),iv)$ are adapted from the proofs of [5, Theorem 17.2, Lemma 17.3.]
Furthermore, since by Lemma 5 iii), a geodesic induced by an optimal $W_{2,Y}$ plan is weakly continuous, we obtain the following proposition from [6, Proposition 8.1.8] which gives a connection to a flow equation and is needed for flow matching.
Proposition 7.
@@
    $\mu_{0,y},\mu_{1,y}$ for $P_{Y}$-a.e. $y\in\mathbb{R}^{d}$ are absolutely continuous with densities $\rho_{0,y},\rho_{1,y}$ which are supported on open, convex, bounded, connected, subsets $\Omega_{0,y},\Omega_{1,y}$ on which they are bounded away from 00 and $\infty$. Assume further that $\rho_{0,y}\in C^{2}(\Omega_{0,y}),\rho_{1,y}\in C^{2}(\Omega_{1,y})$ and let $T_{y}$ be the associated optimal transport maps and $\alpha\in\Gamma^{4}_{Y}(\mu_{0},\mu_{1})$ be the associated optimal transport plan.
@@ 6 Relaxation of the conditional Wasserstein distance
    We cannot use standard optimal transport algorithms like [22] out of the box.
@@
i.e. $\mu_{n}\to\mu$ and $\nu_{n}\to\nu$ as $n\to\infty$ in the $W_{2}$-topology. Let $Y_{n}$ be a random variable distributed like $\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i}}$. Then with probability one $\Gamma_{Y_{n}}^{4}(\mu_{n},\nu_{n})$ contains exactly one plan $\alpha_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},x_{i},y_{i},z_{i}}$ which is clearly optimal. Let $\Delta:\mathbb{R}^{3}\to\mathbb{R}^{4}$ be defined by $\Delta(y,x,z)=(y,x,y,z)$. Then $\hat{\alpha}:=\lim_{n}\alpha_{n}=\Delta_{\sharp}\left(P_{Y}\otimes P_{X}\otimes P_{Z}\right)$ in the $W_{2}$-topology. Moreover, $\hat{\alpha}\in\Gamma^{4}_{Y}(\mu,\nu)$ and
@@
The same idea has been pursued in [2], where the authors rescaled the $y$-costs to obtain a blocktriangular map in the Knothe-Rosenblatt setting [37, 51] and similarly in [31]. Note that [31] was published on ArXiv after our first version [15] of the present paper.
Proposition 10.
@@
Remark 11.
Alternatively, instead of rescaling the costs $d_{\beta}$ we would also rescale the inputs, which was done for instance in [20, 26]. Take for instance (as we do numerically) the cost function $d_{\beta}^{2}=\|x_{1}-x_{2}\|^{2}+\beta\|y_{1}-y_{2}\|^{2}$. Then this is equivalent to rescaling the $Y$-component by $\sqrt{\beta}$.
@@ 7 Bayesian flow matching
Remark 13.
Usually, in conditional generative modelling, the word "conditional" appears the context of inverse problems (or solving class conditional problems). However, in the vanilla flow matching [39] the word "conditional" is used for the loss and the whole procedure is referred to as "conditional flow matching". Therefore, we will call the flow matching procedure for inverse problems "Bayesian flow matching".
Flow matching and OT flow matching aim to sample from a target distribution $P_{X}$ by learning the velocity field $v_{t}$ of a flow ODE [16]
@@
which transports samples from an initial distribution $P_{Z}$ to those from $P_{X}$. Once an approximate velocity field $v^{\theta}_{t}$ is learned, it can be replaced in (26) to get a neural ODE [16].
Flow Matching
The flow matching framework [39, 40] learns $v^{\theta}_{t}$ based on linear interpolation between independent $Z$ and $X$, i.e.
$X_{t}=(1-t)\ Z+t\ X$
and consequently
$X-Z=\frac{d}{dt}X_{t}=v_{t}(X_{t}).$
In a minibatch setting with $(\bm{z},\bm{x})=\left((z_{i},x_{i})\right)_{i=1}^{I}$ sampled from the product distribution $P_{Z}\otimes P_{X}$, this becomes
$v_{t}(\bm{x}_{t})\coloneqq\bm{x}-\bm{z},\quad\bm{x}_{t}:=(1-t)\bm{z}+t\bm{x}.$ (28)
@@
The objective $L_{FM}$ can be also derived differently, with ideas inspired by the score matching framework [32, 57]. Then instead of regressing to the true velocity field at $x_{t}$, they show that regressing to it has the same gradients when one conditions at $x$ ("conditional" flow matching [39, Theorem 2]), which leads to the same loss formulation.
OT Flow Matching
In contrast to the above linear interpolation, the authors in [55, 49] suggested to use the $W_{2}(P_{Z},P_{X})$ coupling $\gamma$, respectively its Monge map $T$ and the corresponding McCann interpolation [43]
@@
Random Bayesian Flow Matching
In [58, 60] flow matching is extended to the conditional setting. Given independent $Z$ and $(Y,X)$ we again consider the linear interpolation between $Z$ and $X$ given by
@@
$(0,X-Z)=\frac{d}{dt}(Y,X_{t})=v_{t}(Y,X_{t}).$
Given a minibatch $(\bm{z},\bm{y},\bm{x})=\left((z_{i},y_{i},x_{i})\right)_{i=1}^{I}$ sampled from the product distribution $P_{Z}\otimes P_{Y,X}$, this becomes
$v_{t}(\bm{y},\bm{x}_{t})\coloneqq(0,\bm{x}-\bm{z}),\quad\bm{x}_{t}:=(1-t)\bm{z}+t\bm{x}.$ (31)
@@
OT Bayesian Flow Matching
For $P_{Y,Z},P_{Y,X}$ as in Proposition 8 there exists an optimal plan $\alpha\in\Gamma_{Y}^{4}(P_{Y,Z},P_{Y,X})$ and corresponding map $T$. Furthermore there exists a vector field $v_{t}\in L^{2}_{\mu_{t}}$ such that there exists a solution $\phi_{t}$ to the flow equation which satisfies
@@ 8 Numerical experiments
In this section, we want to show cases in which it is beneficial to use the conditional Wasserstein distance. First, we verify that the convergence result for an increasing parameter $\beta$ given in Proposition 10 for particle flows to MNIST [19]. Then we show the advantages of our Bayesian OT flow matching procedure on a Gausian mixture model (GMM) toy example and on CIFAR10 [38] class-conditional image generation.
@@ 8.1 Particle flow convergence
In this example, we minimize $W_{Y,d_{\beta}}(P_{Y,X},P_{Y,Z})$ for the empirical measures. We consider the particle flow, i.e., the flow from $(Y,Z)$ to $(Y,X)$ for empirical distributions which minimizes the objective $\mathcal{D}\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},x_{i}},\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},z(t)_{i}}\right)$ for an appropriate distance $\mathcal{D}$, see e.g. [4]. More concretely we follow a particle flow path, i.e., a curve starting with $z(0)_{i}\sim\mathcal{N}(0,I)$ which fulfills
@@
for an appropriate scaling $\eta$ and given joint samples $(y_{i},x_{i})_{i=1}^{n}$. We choose $\mathcal{D}$ as an approximation of $W_{2,d_{\beta}}$ by rescaling $Y$ and using the Sinkorn divergence as the sample based distance measure [23, 21], where the "blur" parameter is chosen so small that it is close to the Wasserstein distance. This way we can numerically verify the convergence in Proposition 10. Note that there are no neural networks involved in this example.
@@ 8.2 GMM example
Here we use an experimental setup from [28]: Recall the Bayesian inverse problem setting in (1), i.e. $Y=f(X)+\Xi$ where choose $P_{X}$ to be a GMM in $\mathbb{R}^{5}$ with 10 mixture components, uniformly chosen means in $[-1,1]$ and standard deviation 0.1. We apply a linear diagonal forward model $f=(f_{i,j})_{i,j=1}^{5}\in\mathbb{R}^{5\times 5}$ with $f_{i,i}=\frac{0.1}{i+1}$ and zero components otherwise and choose $\Xi=\mathcal{N}(0,0.1)$ as a standard Gaussian distribution with standard deviation $0.1$. This yields a posterior measure $P_{X|Y=y}$ which is a also a GMM [28, Lemma 11]. Therefore we can sample and evaluate the true posterior as groundtruth. We train a random Bayesian flow matching model according to $L_{Y,\text{FM}}$(32) and our OT Bayesian flow matching according to $L_{Y,\text{OT}}$(33) with the python package POT [22] on a fixed dataset of size $10000$, where we choose the best model according to a validation set of size $2000$ until the validation loss converges. For both approaches we use the same feed-forward neural network which contains around 140k parameters. We evaluate them using the Sinkhorn distance with blur 0.05 [23] with the package GeomLoss [21] averaged with 100 posteriors and over 3 training and test runs with randomly sampled mixtures. The sampling is done via an explicit Euler discretization of 10 steps. Our proposed OT Bayesian flow matching model trained according to $L_{Y,\text{OT}}$ with $\beta=20$ obtains an average Sinkhorn distance of 0.0217, whereas the random Bayesian flow matching model obtains a value of 0.0278. In Figure 2 one can see that both models approximate the posterior very well.
Figure 2: Posterior histograms for different methods with random Bayesian flow matching on the left and our OT Bayesian flow matching on the right. Ground truth posterior is in orange and model prediction in blue.

### v3

@@ 5 Geodesics and Velocity Fields
In this section, we deal with geodesics and velocity fields in $\left(\mathcal{P}_{Y,2}(\mathbb{R}^{d}\times\mathbb{R}^{m}),W_{2,Y}\right)$. We restrict our attention to $p=2$ and $A=\mathbb{R}^{d}$, $B=\mathbb{R}^{m}$. Coming from inverse problems, we have considered probability measures $P_{Y,X}$ related to random variables $(Y,X)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$. When switching to flows, it is more convenient to address equivalently just probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{m}$.
Let us recall some results, which can be found, e.g. in (Ambrosio et al., 2005) for our setting. A curve $\mu\colon[0,1]\to\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ is called a geodesic if
@@
The Wasserstein space is geodesic, i.e. any two measures $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ can be connected by a geodesic. Let $e_{t}:(\mathbb{R}^{d}\times\mathbb{R}^{m})^{2}\to\mathbb{R}^{n}\times\mathbb{R}^{m}$, $t\in[0,1]$ be defined by
@@
Conversely, any optimal plan $\alpha\in\Gamma(\mu_{0},\mu_{1})$ gives by (16) rise to a geodesic connecting $\mu_{0}$ and $\mu_{1}$. The following lemma considers curves defined by (16) which connect measures $\mu_{0},\mu_{1}\in\mathcal{P}_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$. Its proof is given in the appendix and is similar to (Ambrosio et al., 2005, Theorem 7.2.2).
Lemma 5
@@
By the following proposition, the above geodesic $\mu_{t}$ has an associated vector field $v_{t}$ such that $(\mu_{t},v_{t})$ satisfy a continuity equation. Moreover, informally speaking, the associated vector field $v_{t}$ does not transport any mass in the $y$-component. This is related to the observation in (Du et al., 2023, Section 4.2) and (Peszek and Poyato, 2023, Proposition 3.21).
Proposition 6
@@
The proof is given in Appendix D and parts $i),ii),iv)$ are adapted from the proofs of (Ambrosio et al., 2021, Theorem 17.2, Lemma 17.3.)
Furthermore, since by Lemma 5 iii), a geodesic induced by an optimal $W_{2,Y}$ plan is weakly continuous, we obtain the following proposition from (Ambrosio et al., 2005, Proposition 8.1.8) which gives a connection to a flow equation and is needed for flow matching.
Proposition 7
@@
    $\mu_{0,y},\mu_{1,y}$ for $P_{Y}$-a.e. $y\in\mathbb{R}^{d}$ are absolutely continuous with densities $\rho_{0,y},\rho_{1,y}$ which are supported on open, convex, bounded, connected, subsets $\Omega_{0,y},\Omega_{1,y}$ on which they are bounded away from 0 and $\infty$. Assume further that $\rho_{0,y}\in C^{2}(\Omega_{0,y}),\rho_{1,y}\in C^{2}(\Omega_{1,y})$ and let $T_{y}$ be the associated optimal transport maps and $\alpha\in\Gamma^{4}_{Y}(\mu_{0},\mu_{1})$ be the associated optimal transport plan.
@@
The following proposition is a consequence of (González-Sanz and Sheng, 2024, Corollary 1.2). We only give an informal formulation here and refer for the details to Proposition 17 in the appendix. Notably, this helps us to overcome measurability issues when working with continuous $P_{Y}$ and therefore is applicable to a broader class of inverse problems.
Proposition 9
Let $P_{Y}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, $\mu_{0}=P_{Y}\times\mu^{Z}_{0}$ and let $\mu_{1}=\mu_{1}^{y}\times_{y}P_{Y}$ with density $p_{1}^{y}$ of $\mu_{1}^{y}$. Assume further that the map $y\mapsto p_{1}^{y}$ is a $C^{1}$ map. Under relatively mild assumptions, see Proposition 17, the following statements hold true. There exists a $W_{2,Y}$-optimal transport map $T:(y,x)\mapsto(y,T_{y}(x))$ i.e. $\alpha=(\operatorname{Id},T)_{\sharp}\mu_{0}\in\Gamma_{o,Y}(\mu_{0},\mu_{1})$, where $T_{y}$ is the optimal transport map for $\mu_{0}^{Z}$ and $\mu_{1}^{y}$. Let $\mu_{t}\coloneqq(e_{t})_{\sharp}\alpha$ with associated vector field $v_{t}\in L^{2}_{\mu_{t}}$, where $(v_{t})_{j}=0$ for all $j\leq d$. Then there is a representative of $v_{t}$ such that the flow equation
$\frac{\mathrm{d}}{\mathrm{d}t}\phi_{t}=v_{t}(\phi_{t});\quad\phi_{0}(y,x)=(y,x)$ (24)
admits a global solution and $\mu_{t}=(\phi_{t})_{\sharp}\mu_{0}$. Furthermore, we have
$v_{t}(\phi_{t}(y,x))=T(y,x)-(y,x)=\left(0,T_{y}(x)-x\right)$ (25)
for $\mu_{0}$-a.e. $(y,x)\in\mathbb{R}^{d}\times\mathbb{R}^{m}$.
@@ 6 Relaxation of the Conditional Wasserstein Distance
    We cannot use standard optimal transport algorithms like (Flamary et al., 2021) out of the box.
@@
i.e. $\mu_{n}\to\mu$ and $\nu_{n}\to\nu$ as $n\to\infty$ in the $W_{2}$-topology. Note that we cannot compare $\mu_{n}$ and $\mu$ in the $W_{2,Y}$ topology since $\mu_{n}\notin\mathcal{P}_{2,Y}(\mathbb{R}\times\mathbb{R})$. Let $Y_{n}$ be a random variable distributed like $\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i}}$. Then with probability one $\Gamma_{Y_{n}}^{4}(\mu_{n},\nu_{n})$ contains exactly one plan $\alpha_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},x_{i},y_{i},z_{i}}$ which is clearly optimal. Let $\Delta:\mathbb{R}^{3}\to\mathbb{R}^{4}$ be defined by $\Delta(y,x,z)=(y,x,y,z)$. Then $\hat{\alpha}:=\lim_{n}\alpha_{n}=\Delta_{\sharp}\left(P_{Y}\otimes P_{X}\otimes P_{Z}\right)$ in the $W_{2}$-topology. Moreover, $\hat{\alpha}\in\Gamma^{4}_{Y}(\mu,\nu)$ and
@@
The same idea has been pursued in (Alfonso et al., 2023), where the authors rescaled the $y$-costs to obtain a block-triangular map in the Knothe-Rosenblatt setting (Knothe, 1957; Rosenblatt, 1952) and similarly in (Hosseini et al., 2024). Note that (Hosseini et al., 2024) was published on ArXiv after our first version (Chemseddine et al., 2023) of the present paper. The distance $W_{2,\beta}$ was also discussed in (Peszek and Poyato, 2023, Proposition 3.10) and Proposition 10 can be found in the proof thereof.
Proposition 10
@@
Remark 11
Alternatively, instead of rescaling the costs $d_{\beta}$ we would also rescale the inputs, which was done for instance in (Du et al., 2023; Hagemann et al., 2024). Take for instance (as we do numerically) the cost function $d_{\beta}^{2}=\|x_{1}-x_{2}\|^{2}+\beta\|y_{1}-y_{2}\|^{2}$. Then this is equivalent to rescaling the $Y$-component by $\sqrt{\beta}$.
@@ 7 Bayesian Flow Matching
Remark 13
Usually, in conditional generative modeling, the word "conditional" appears the context of inverse problems (or solving class conditional problems). However, in the vanilla flow matching (Lipman et al., 2023) the word "conditional" is used for paths fixing a target sample and the whole procedure is referred to as "conditional flow matching". Therefore, we will call the flow matching procedure for inverse problems "Bayesian flow matching".
Flow matching and OT flow matching aim to sample from a target distribution $P_{X}$ by learning the velocity field $v_{t}$ of a flow ODE (Chen et al., 2018)
@@
which transports samples from an initial distribution $P_{Z}$ to those from $P_{X}$. Once an approximate velocity field $v^{\theta}_{t}$ is learned, it can be replaced in (28) to get a neural ODE (Chen et al., 2018).
Flow Matching The flow matching framework (Lipman et al., 2023; Liu et al., 2023) learns $v^{\theta}_{t}$ based on linear interpolation between independent $Z$ and $X$, i.e.
$X_{t}=(1-t)\ Z+t\ X,$
and in (Liu et al., 2023, Theorem D.3) it was shown that a suitable vector field is
$v_{t}(x)=\mathbb{E}\left[\frac{d}{dt}X_{t}|X_{t}=x\right]=\mathbb{E}[X-Z|X_{t}=x].$
@@
The objective $L_{FM}$ can be also derived differently, with ideas inspired by the score matching framework (Hyvärinen and Dayan, 2005; Vincent, 2011). Then, instead of regressing to the true velocity field at $x_{t}$, they show that regressing to it has the same gradients when one conditions at $x$ ("conditional" flow matching (Lipman et al., 2023, Theorem 2)), which leads to the same loss formulation.
OT Flow Matching In contrast to the above linear interpolation, the authors in (Tong et al., 2023; Pooladian et al., 2023) suggested to use the $W_{2}(P_{Z},P_{X})$ coupling $\gamma$, respectively its Monge map $T$ and the corresponding McCann interpolation (McCann, 1997)
@@
Random Bayesian Flow Matching In (Wildberger et al., 2023; Zheng et al., 2023) flow matching is extended to the conditional setting. Given independent $Z$ and $(Y,X)$ we again consider the linear interpolation between $Z$ and $X$ given by
@@
$(0,X-Z)=\frac{d}{dt}(Y,X_{t}).$
@@
OT Bayesian Flow Matching For $P_{Y,Z},P_{Y,X}$ as in Proposition 8 or Proposition 9 there exists an optimal plan $\alpha\in\Gamma_{Y}^{4}(P_{Y,Z},P_{Y,X})$ and corresponding map $T$. Furthermore there exists a vector field $v_{t}\in L^{2}_{\mu_{t}}$ such that there exists a solution $\phi_{t}$ to the flow equation which satisfies
@@ 8 Numerical Experiments
In this section, we want to show cases in which it is beneficial to use the conditional Wasserstein distance. First, we verify that the convergence result for an increasing parameter $\beta$ given in Proposition 10 for particle flows to MNIST (Deng, 2012). Then we show the advantages of our Bayesian OT flow matching procedure on a Gausian mixture model (GMM) toy example and on CIFAR10 (Krizhevsky et al., 2009) class-conditional image generation.
@@ 8.1 Particle Flow Convergence
In this example, we minimize $W_{Y,d_{\beta}}(P_{Y,X},P_{Y,Z})$ for the empirical measures. We consider the particle flow, i.e., the flow from $(Y,Z)$ to $(Y,X)$ for empirical distributions which minimizes the objective $\mathcal{D}\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},x_{i}},\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},z(t)_{i}}\right)$ for an appropriate distance $\mathcal{D}$, see e.g. (Altekrüger et al., 2023). More concretely we follow a particle flow path, i.e., a curve starting with $z(0)_{i}\sim\mathcal{N}(0,I)$ which fulfills
@@
for an appropriate scaling $\eta$ and given joint samples $(y_{i},x_{i})_{i=1}^{n}$. We choose $\mathcal{D}$ as an approximation of $W_{2,d_{\beta}}$ by rescaling $Y$ and using the Sinkorn divergence as the sample based distance measure (Genevay et al., 2018; Feydy et al., 2019), where the "blur" parameter is chosen so small that it is close to the Wasserstein distance. This way we can numerically verify the convergence in Proposition 10. Note that there are no neural networks involved in this example.
@@ 8.2 GMM Example
Here we use an experimental setup from (Hagemann et al., 2022b): Recall the Bayesian inverse problem setting in (1), i.e. $Y=f(X)+\Xi$ where we choose $P_{X}$ to be a GMM in $\mathbb{R}^{5}$ with 10 mixture components, uniformly chosen means in [1,1][-1,1] and standard deviation 0.1. We apply a linear diagonal forward model $f=(f_{i,j})_{i,j=1}^{5}\in\mathbb{R}^{5\times 5}$ with $f_{i,i}=\frac{0.1}{i+1}$ and zero components otherwise and choose $\Xi=\mathcal{N}(0,0.1)$ as a standard Gaussian distribution with standard deviation $0.1$. This yields a posterior measure $P_{X|Y=y}$ which is a also a GMM (Hagemann et al., 2022b, Lemma 11). Therefore we can sample and evaluate the true posterior as groundtruth. We train a random Bayesian flow matching model according to $L_{Y,\text{FM}}$(32) and our OT Bayesian flow matching according to $L_{Y,\text{OT}}$(33) with the python package POT (Flamary et al., 2021) on a fixed dataset of size $10000$, where we choose the best model according to a validation set of size $2000$ until the validation loss converges. For both approaches we use the same feed-forward neural network which contains around 140k parameters. We evaluate them using the Sinkhorn distance with blur 0.05 (Genevay et al., 2018) with the package GeomLoss (Feydy et al., 2019) averaged with 100 posteriors and over 10 training and test runs with randomly sampled mixtures. The sampling is done via an explicit Euler discretization of 10 steps. Our proposed OT Bayesian flow matching model trained according to $L_{Y,\text{OT}}$ with $\beta=20$ obtains an average Sinkhorn distance of 0.0225, whereas the random Bayesian flow matching model obtains a value of 0.0247. In Figure 2 one can see that both models approximate the posterior very well. We repeated the same experiments but used only $3$ Euler steps for sampling. In this case we obtained an average Sinkhorn distance of 0.0760 for the OT Bayesian flow matching and a value of 0.1044 for the random Bayesian flow matching. This indicates that the paths learned by our method are more straight.
Figure 2: Posterior histograms for different methods with random Bayesian flow matching on the left and our OT Bayesian flow matching on the right for 10 Euler steps. Ground truth posterior is in orange and model prediction in blue.